OpenStudy (anonymous):

What is the solution of the system? Use substitution. x = –2y + 10 –6y = 6x – 6 z = –3x + 10y A. x = 9, y = –8, z = 114 B. x = –8, y = 9, z = 114 C. x = –8, y = 114, z = 9 D. x = 114, y = 9, z = –8 A, B, C, or D? Please explain (:

OpenStudy (anonymous):

use the first equation, the x value is now -2y+10 and use that in the second equation. solve for a and y using just those 2 equations. then plug in the values of x and y in the 3rd eqtion and get z. sorry I can't do it all now. I have 2 exams tomorrow, gotta study more and go to sleep. been studying for almost 8 hours now. bye

OpenStudy (anonymous):

x = –2y + 10 –6y = 6x – 6 z = –3x + 10y 2nd line yields: y=-x+1=2y-10+1=2y-9 y=9 x=-18+10=-8 z=-3(-8)+10(9) 24+90=114 answer is B now bye

OpenStudy (anonymous):

alright, thanks a lot. good night (: